OpenStudy (knespinoza):

Imagine yourself diving off a platform 10 m above the water below. a. If you walk very slowly off the platform, so that you fall directly downward, roughly how long will it take for you to reach the water?

OpenStudy (osprey):

You fall with the acceleration of gravity (I wouldn't try it from 10m unless I knew what I was doing). Initial speed is zero. s=1/2 a t squared. t = square root (2s/a) seconds which is sqrt(2x10/10) = about 1.414 seconds. Mind you it would probably feel like a lifetime if you hadn't prepared for it. DO NOT try this, please. Bon chance, et bon voyage. http://perendis.webs.com

OpenStudy (irishboy123):

Drawing it always helps. And horizontal motion shouldn't affect the answer. In our idealised world.

OpenStudy (osprey):

@IrishBoy123 That's what i did in http://perendis.webs.com, albeit a pretty lousy and illegible drawing which is only half relevant here ... Where IS this idealised world ? Never Never land ?

OpenStudy (mathmale):

The governing equation here is\[s=s _{0}+v _{0}t+\frac{ 1 }{ 2 }at^2\] Have you seen this before? Think: What do s_0, v_0, a and t represent?